A little car of mass M1 has a maximum acceleration of 2.57m/s^2 when applying a force F1. What is the new maximum acceleration of the little car if it tows another car that has the same mass M1 and the same applied force F1?

Grade:12

1 Answers

Vikas TU
5 years ago
For little car without towing,
F1 = M1*2.57.........................(1)
When towing another car of same mass M1, with same force F1:
F1 = (M1 + M1)*a
F1 = 2M1*a
a = F1/2M1
from eqn (1) put the value we get,
a = 2.57M1/2M1 => 2.57/2 => 1.285 m/s^2